# MolCRAFT: Structure-Based Drug Design in Continuous Parameter Space

### v2

@@ Abstract
Generative models for structure-based drug design (SBDD) have shown promising results in recent years. Existing works mainly focus on how to generate molecules with higher binding affinity, ignoring the feasibility prerequisites for generated 3D poses and resulting in false positives. We conduct thorough studies on key factors of ill-conformational problems when applying autoregressive methods and diffusion to SBDD, including mode collapse and hybrid continuous-discrete space. In this paper, we introduce MolCRAFT, the first SBDD model that operates in the continuous parameter space, together with a novel noise reduced sampling strategy. Empirical results show that our model consistently achieves superior performance in binding affinity with more stable 3D structure, demonstrating our ability to accurately model interatomic interactions. To our best knowledge, MolCRAFT is the first to achieve reference-level Vina Scores (-6.59 kcal/mol) with comparable molecular size, outperforming other strong baselines by a wide margin (-0.84 kcal/mol).
@@ 1 Introduction
Structure-based drug design (SBDD) advances drug discovery by leveraging 3D structures of biological targets, thereby facilitating efficient and rational design of molecules within a certain chemical space of interests (Wang et al., 2022; Isert et al., 2023). In recent years, the generative model for molecules has emerged as a promising direction, which could streamline SBDD by directly proposing desired molecules, eliminating the need for exhaustive blind search in the vast space (Walters, 2019; Luo et al., 2021). Recent progress in SBDD can be divided into two categories, i.e. auto-regressive models (Luo et al., 2021; Peng et al., 2022; Zhang et al., 2023) as next-token prediction for text generation, and diffusion models (Guan et al., 2022, 2023) as for image generation.
The essential criteria for drug-like candidate molecules are outlined as follows: (i) high affinity towards specific binding sites (a.k.a, protein pockets), where a higher affinity indicates better performance, (ii) satisfactory drug-like properties, such as synthesizability and drug-likeness scores, which often serve as thresholds for filtering out unfavorable compounds (Ursu et al., 2011; Tian et al., 2015), and (iii) well-conformational 3D structure, which needs special attention for SBDD models, because they risk generating unrealistic molecular 3D conformations yet with deceptively high affinities.
@@
Figure 1: Typical resulting implausible molecules from generative models. (a) Unusual 3-membered rings generated by AR, large fused rings with more than 7 atoms generated by diffusion models. (b) Examples of steric clashes by FLAG, and other ligand undergoing significant conformational rearrangements upon redocking (Before: blue. After: green). (c) Failures in generation process. Left: atoms mis-connected in autoregressive sampling. Right: incomplete molecules with multiple components.
In order to capture the complicated data manifold for molecules, we take a shift to a unified continuous parameter space instead of a hybrid space, inspired by Graves et al. (2023). We propose MolCRAFT (Continuous paRAmeter space Facilitated molecular generaTion), which not only alleviates the mode collapse issue by non-autoregressive generation as in its diffusion counterparts, but also addresses the continuous-discrete gap by applying continuous noise and smooth transformation, leading to high-affinity as well as well-conformational drug candidates.
@@
  • We investigate the false positive phenomenons of current SBDD models, and identify several key problems including the mode collapse of autoregressive methods, and the gap of continuous-discrete space when applying diffusion models.
@@
  • We conduct comprehensive evaluation under controlled molecular sizes. Experiments show that our model generates high-affinity binders with feasible 3D poses. To our best knowledge, we are the first to achieve reference-level Vina Scores (-6.59 kcal/mol, compared to reference -6.36 kcal/mol) with comparable molecule size, outperforming other strong baselines by a wide margin (-0.84 kcal/mol).
@@ 2.1 Failure Modes of Generated Molecules
    Distorted geometry. We visualize the generated molecules at median strain energy (see Table 2), and models tend to produce either too many uncommon 3- or 4-member rings, or extra-large rings with unstable structures, leading to much higher strain energy.
@@ 2.3 Hybrid Continuous-Discrete Space
To further illustrate the difference between continuous-discrete diffusion and our fully continuous MolCRAFT, we sample 10 molecules for each of the 100 test proteins, and plot the curves of the ratio of valid molecules, complete molecules against different timesteps during sampling. As shown in Fig. 3, continuous-discrete diffusions heavily rely on the latter steps, passing a certain validity and completeness threshold in the final 60%-90% stage where noise scales are lower, while MolCRAFT approaches target distribution far earlier (in the first 20%-40% steps), thereby possessing greater capacity to progressively refine and adjust the generated feasible structures, resulting in better conformations.
@@ 4.1 Resolving Different Modalities in Parameter Space
where $\mathbf{e}_{\mathbf{v}_{M}}=\Big{[}\mathbf{e}_{\mathbf{v}_{M}^{(1)}},\dots,\mathbf{e}_{\mathbf{v}_{M}^{(K)}}\Big{]}\in\mathbb{R}^{N_{M}\times K}$, $\mathbf{e}_{j}\in\mathbb{R}^{K}$ is the projection from the class index $j$ to the length-$K$ one-hot vector, and $K$ the number of atom types. Note that we could set different noise schedules for different modalities ($\alpha$ for coordinates and $\alpha^{\prime}$ for types) for more efficient training of the joint noise prediction network.
@@
Denote the SE-(3) transformation as $T_{g}$, the likelihood is invariant w.r.t. $T_{g}$ on the protein-molecule complex: $p_{\phi}(T_{g}(\mathbf{m}|\mathbf{p}))=p_{\phi}(\mathbf{m}|\mathbf{p})$ if we shift the Center of Mass (CoM) of protein atoms to zero and parameterize the output network $\boldsymbol{\Phi}(\boldsymbol{\theta},\mathbf{p},t)$ with an SE-(3) equivariant network.
@@ 5 Experiments
Table 2: Summary of different properties of reference and generated molecules under different sizes. (↑) / (↓) indicates larger / smaller is better. Top 2 results are highlighted with bold text and underlined text. Note: SE is short for Strain Energy, Div for Diversity, BF for Binding Feasibility, and SR for Success Rate.
@@
Reference -6.36 -6.46 -6.71 -6.49 -7.45 -7.26 38 198 5.57 34.0 0.73 0.48 - 26.0 25.0 22.8
AR -5.75 -5.64 -6.18 -5.88 -6.75 -6.62 260 2287 4.36 36.5 0.63 0.51 0.70 16.1 6.9 17.7
Pocket2Mol -5.14 -4.70 -6.42 -5.82 -7.15 -6.79 102 373 6.10 32.0 0.76 0.57 0.69 23.8 24.4 17.7
FLAG 45.85 36.52 9.71 -2.43 -4.84 -5.56 25 4384 68.55 0.3 0.63 0.61 0.70 0.0 1.8 16.7
Ours-small -5.96 -5.89 -6.34 -6.04 -6.98 -6.63 44 275 4.77 39.5 0.74 0.52 0.74 33.3 17.4 17.8
TargetDiff -5.47 -6.30 -6.64 -6.83 -7.80 -7.91 368 13527 11.13 37.1 0.58 0.48 0.72 13.5 10.5 24.2
Decomp-R -5.19 -5.27 -6.03 -6.00 -7.03 -7.16 111 1217 7.92 24.2 0.66 0.51 0.73 14.6 14.9 21.2
Ours -6.59 -7.05 -7.24 -7.26 -7.80 -7.92 84 517 7.02 46.1 0.69 0.50 0.72 35.9 26.0 22.7
Decomp-O -5.67 -6.04 -7.04 -7.09 -8.39 -8.43 368 3876 13.76 27.2 0.61 0.45 0.68 11.1 24.5 29.4
Ours-large -6.61 -8.16 -8.14 -8.45 -9.21 -9.22 174 1079 10.87 45.0 0.62 0.46 0.61 31.1 36.6 29.4
@@ 5.2 Main Results
I. Our model consistently outperforms other strong baselines in affinities, achieving a reference-level Vina Score of -6.59 kcal/mol. As Vina Score directly scores the pose and Vina Min only optimizes locally, they directly measure the generated pose quality. To the best of our knowledge, MolCRAFT is the first to achieve reference-level affinity scores without significant rearrangements via redocking, which demonstrates our superiority in learning binding interactions for SBDD.
@@
IV. Our binding complex contains fewer clashes and remains consistent after redocking. We achieve few steric clashes, and has the best RMSD performance, which means 46% of our molecules already resemble accurate docking pose even without force field optimization or redocking, rendering it reliable for generating molecules in 3D space. The reason why we achieve even better RMSD than reference could be explained by a distribution shift from the training set to the reference set. In the construction of dataset [^2] There are two kinds of 3D ligand poses in the dataset, i.e. Vina minimized poses in the given receptor, and Vina docked poses. https://github.com/gnina/models/tree/master/data/CrossDocked2020, the training set contains 52.4% docked molecules, while the test set only contains 37.0% docked ones, which aligns with the phenomenon that there are only 34.0% reference molecules with RMSD $<$ 2Å upon redocking. This accounts for why MolCRAFT has more consistent and high-affinity binders, which effectively captures the training set distribution and learns the binding dynamics.
@@ 7 Conclusion
Broader Impact
@@ D.1 Full Evaluation Results
In addition to our main results in Table 2, we provide evaluation in terms of key interactions in Table 4, i.e., No. Hydrogen Bond Donors (HB Donors), No. Hydrogen Bond Acceptors (HB Acceptors), van-der Waals contacts (vdWs) and Hydrophobic interactions as described in PoseCheck (Harris et al., 2023). It can be seen that, under different molecule sizes, ours is the best in forming hydrogen bonds, indicating that MolCRAFT finely captures key protein-ligand interactions.

### v4

@@ Abstract
Generative models for structure-based drug design (SBDD) have shown promising results in recent years. Existing works mainly focus on how to generate molecules with higher binding affinity, ignoring the feasibility prerequisites for generated 3D poses and resulting in false positives. We conduct thorough studies on key factors of ill-conformational problems when applying autoregressive methods and diffusion to SBDD, including mode collapse and hybrid continuous-discrete space. In this paper, we introduce MolCRAFT, the first SBDD model that operates in the continuous parameter space, together with a novel noise reduced sampling strategy. Empirical results show that our model consistently achieves superior performance in binding affinity with more stable 3D structure, demonstrating our ability to accurately model interatomic interactions. To our best knowledge, MolCRAFT is the first to achieve reference-level Vina Scores (-6.59 kcal/mol) with comparable molecular size, outperforming other strong baselines by a wide margin (-0.84 kcal/mol). Code is available at https://github.com/AlgoMole/MolCRAFT.
@@ 1 Introduction
Structure-based drug design (SBDD) advances drug discovery by leveraging 3D structures of biological targets, thereby facilitating efficient and rational design of molecules within a certain chemical space of interests (Wang et al., 2022; Isert et al., 2023). Recently, generative models for molecules have emerged as a promising direction, which could streamline SBDD by directly proposing desired molecules, eliminating the need for exhaustive blind search in the vast space (Walters, 2019; Luo et al., 2021). Recent progress in SBDD can be divided into two categories, i.e. auto-regressive models (Luo et al., 2021; Peng et al., 2022; Zhang et al., 2023) as next-token prediction for text generation, and diffusion models (Guan et al., 2022, 2023) as for image generation.
The essential criteria for drug-like candidate molecules are outlined as follows: (i) high affinity towards specific binding sites (a.k.a, protein pockets), where a higher affinity indicates better performance, (ii) satisfactory drug-like properties, such as synthesizability and drug-likeness scores, which often serve as thresholds for filtering out unfavorable compounds (Ursu et al., 2011; Tian et al., 2015), and (iii) well-conformational 3D structure, which needs special attention as there is a danger of generating unrealistic 3D conformations yet with deceptively high affinities.
@@
Figure 1: Typical failure modes. (a) Unusual 3-membered rings generated by AR, large fused rings with more than 7 atoms generated by diffusion models. (b) Examples of steric clashes by FLAG, and other ligand undergoing significant conformational rearrangements upon redocking (Before: blue. After: green). (c) Failures in generation process. Left: atoms mis-connected in autoregressive sampling. Right: incomplete molecules with multiple components.
In order to capture the complicated data manifold for molecules, we take a shift to a unified continuous parameter space instead of a hybrid space, inspired by Graves et al. (2023). We propose MolCRAFT (Continuous paRAmeter space Facilitated molecular generaTion), which not only alleviates the mode collapse issue by non-autoregressive generation as in its diffusion counterparts, but also addresses the continuous-discrete gap by applying continuous noise and smooth transformation.
@@
  • We investigate the challenges of current SBDD models, and identify key issues including the mode collapse of autoregressive methods, and the gap of continuous-discrete space when applying diffusion models.
@@
  • We conduct comprehensive evaluation under controlled molecular sizes. Experiments show that our model generates high-affinity binders with feasible 3D poses. To our best knowledge, we are the first to achieve reference-level Vina Scores (-6.59 vs. -6.36 kcal/mol) with comparable molecule size, outperforming other strong baselines by a wide margin (-0.84 kcal/mol).
@@ 2.1 Failure Modes of Generated Molecules
    Distorted geometry. We visualize the generated molecules at median strain energy (see Table 2), and models tend to produce either too many uncommon 3- or 4-membered rings, or extra-large rings with unstable structures, leading to much higher strain energy.
@@ 2.3 Hybrid Continuous-Discrete Space
To further illustrate the difference between continuous-discrete diffusion and our fully continuous MolCRAFT, we sample 10 molecules for each of the 100 test proteins, and plot the curves of the ratio of valid, complete molecules against different timesteps during sampling. As shown in Fig. 3, continuous-discrete diffusions heavily rely on the latter steps, passing a certain validity and completeness threshold in the final 60%-90% stage where noise scales are lower, while MolCRAFT approaches target distribution far earlier (in the first 20%-40% steps), thereby possessing greater capacity to progressively refine and adjust the generated feasible structures, resulting in better conformations.
@@ 4.1 Resolving Different Modalities in Parameter Space
where $\mathbf{e}_{\mathbf{v}_{M}}=[\mathbf{e}_{\mathbf{v}_{M}^{(1)}},\dots,\mathbf{e}_{\mathbf{v}_{M}^{(N_{M})}}]\in\mathbb{R}^{N_{M}\times K}$, $\mathbf{e}_{j}\in\mathbb{R}^{K}$ is the projection from the class index $j$ to the length-$K$ one-hot vector, and $K$ the number of atom types. Note that we could set different noise schedules for different modalities ($\alpha$ for coordinates and $\alpha^{\prime}$ for types) for more efficient training of the joint noise prediction network.
@@
Denote the SE-(3) transformation as $T_{g}$, the likelihood is invariant w.r.t. $T_{g}$ on the protein-ligand complex: $p_{\phi}(T_{g}(\mathbf{m}|\mathbf{p}))=p_{\phi}(\mathbf{m}|\mathbf{p})$ if we shift the Center of Mass (CoM) of protein atoms to zero and parameterize $\boldsymbol{\Phi}(\boldsymbol{\theta},\mathbf{p},t)$ with an SE-(3) equivariant network.
@@ 5 Experiments
Table 2: Summary of different properties of reference and generated molecules under different sizes. (↑) / (↓) indicates larger / smaller is better. Top 2 results are highlighted with bold text and underlined text. Note: SE is short for Strain Energy, Div for Diversity, BF for Binding Feasibility, and SR for Success Rate. Baselines are either evaluated based on publicly available codebase (Decomp-R) or officially released samples (others).
@@
Reference -6.36 -6.46 -6.71 -6.49 -7.45 -7.26 34 107 196 5.51 34.0 0.73 0.48 - 29.0 25.0 22.8
AR -5.75 -5.64 -6.18 -5.88 -6.75 -6.62 259 595 2286 4.49 31.1 0.64 0.51 0.70 17.3 7.1 17.7
Pocket2Mol -5.14 -4.70 -6.42 -5.82 -7.15 -6.79 102 189 374 6.24 30.8 0.76 0.57 0.69 24.6 24.4 17.7
Ours-small -5.96 -5.89 -6.33 -6.04 -6.98 -6.63 44 103 274 4.88 38.6 0.74 0.52 0.74 33.3 17.4 17.8
FLAG [^2] Though relatively worse than reported in FLAG paper, the evaluation is based on released samples and confirmed by authors. 16.48 4.53 1.21 -4.04 -5.63 -6.61 143 396 1164 40.83 8.2 0.70 0.49 0.70 3.8 14.1 21.5
TargetDiff -5.47 -6.30 -6.64 -6.83 -7.80 -7.91 369 1243 13871 10.84 29.4 0.58 0.48 0.72 14.4 10.5 24.2
Decomp-R -5.19 -5.27 -6.03 -6.00 -7.03 -7.16 115 421 1424 8.16 22.7 0.66 0.51 0.73 14.6 14.9 21.2
Ours -6.59 -7.04 -7.27 -7.26 -7.92 -8.01 83 195 510 7.09 41.8 0.69 0.50 0.72 33.8 26.8 22.7
Decomp-O -5.67 -6.04 -7.04 -7.09 -8.39 -8.43 379 983 4133 14.63 23.9 0.61 0.45 0.68 11.1 24.5 29.4
Ours-large -6.61 -8.14 -8.14 -8.42 -9.25 -9.20 171 333 1110 10.73 42.7 0.62 0.46 0.61 31.1 36.1 29.4
@@ 5.2 Main Results
I. Our model consistently outperforms other strong baselines in affinities, achieving a reference-level Vina Score of -6.59 kcal/mol. As Vina Score directly scores the pose and Vina Min only optimizes locally, they directly measure the generated pose quality. To the best of our knowledge, MolCRAFT is the first to achieve reference-level affinity scores without significant rearrangements via redocking, showing our superiority in learning binding interactions.
@@
IV. Our binding complex contains fewer clashes and remains consistent after redocking. We achieve few steric clashes, and has the best RMSD performance, which means 46% of our molecules already resemble accurate docking pose even without force field optimization or redocking, rendering it reliable for generating molecules in 3D space. The reason why we achieve even better RMSD than reference can be explained by a distribution shift. In dataset construction [^3] There are two kinds of 3D ligand poses in the dataset, i.e. Vina minimized poses in the given receptor, and Vina docked poses. https://github.com/gnina/models/tree/master/data/CrossDocked2020, the training set contains 52.4% docked molecules, while the test set only contains 37.0% docked ones, which aligns with the fact that for reference there are only 34.0% with RMSD $<$ 2Å. This accounts for why MolCRAFT has more consistent and high-affinity binders, which effectively captures the training set distribution and learns the binding dynamics.
@@ 7 Conclusion
Acknowledgements
This work is supported by the National Science and Technology Major Project (2022ZD0117502), Natural Science Foundation of China (62376133) and Guoqiang Research Institute General Project, Tsinghua University (No. 2021GQG1012). The authors would like to thank Xiangyu Li for his valuable comments on this work.
Impact Statement
@@ D.1 Full Evaluation Results
In addition to our main results in Table 2, we provide evaluation in terms of key interactions in Table 4, i.e., No. Hydrogen Bond Donors (HB Donors), No. Hydrogen Bond Acceptors (HB Acceptors), van-der Waals contacts (vdWs) and Hydrophobic interactions as described in PoseCheck (Harris et al., 2023). It can be seen that under different molecule sizes, ours is consistently the best in forming hydrogen bonds, indicating that MolCRAFT finely captures key protein-ligand interactions.